### v1

@@ Abstract
Urban tree biodiversity is critical for climate resilience, ecological stability, and livability in cities, yet most municipalities lack detailed knowledge of their canopies. Field-based inventories provide reliable estimates of Shannon and Simpson diversity but are costly and time-consuming, while supervised AI methods require labeled data that often fail to generalize across regions. We introduce an unsupervised clustering framework that integrates visual embeddings from street-level imagery with spatial planting patterns to estimate biodiversity without labels. Applied to eight North American cities (64,000–560,000 trees, 35–202 genera), the method recovers genus-level diversity patterns with high fidelity, achieving low Wasserstein distances to ground truth for Shannon and Simpson indices and preserving spatial autocorrelation. This scalable, fine-grained approach enables biodiversity mapping in cities lacking detailed inventories and offers a pathway for continuous, low-cost monitoring to support equitable access to greenery and adaptive management of urban ecosystems.
@@
Biodiversity, Unsupervised Learning, Streetviews, Urban Resilience
@@ 1 Introduction
Open access data sources, like Street View Imagery (SVI) offer a scalable alternative for biodiversity monitoring due to their extensive spatial and temporal coverage [22]. However, SVI datasets are unlabeled, meaning the images do not specify the tree species present, and therefore require substantial pre-processing to extract ecological information suitable for Shannon and Simpson index estimation. Identifying individual taxa from imagery is a fine-grained visual recognition problem, characterized by high intra-class variability (e.g., seasonality, age, background) and low inter-class variability between morphologically similar taxa. Current supervised ecological AI models like [23, 24], while effective in their training domains, often under-perform when applied to regions or taxa not well represented in their training data due to geographic domain shift [25].
Self-supervised learning offers a promising alternative by eliminating the need for labeled data [26]. However, when applied to limited or homogeneous datasets, these methods often under-perform due to overfitting and poor generalization, as robust representations typically require large and diverse training samples. However, few clustering approaches have been explicitly designed to preserve abundance distributions in a way that enables accurate Shannon and Simpson estimation from unlabeled imagery. This capability is essential for producing ecologically meaningful biodiversity metrics from large-scale, unlabeled datasets.
In this work, we address the challenge of large-scale, label-free biodiversity monitoring by developing an unsupervised clustering approach that estimates genus-level urban tree diversity directly from street-level imagery. We apply the framework across eight North American cities with contrasting forest compositions and imbalance, and evaluate its ability to recover biodiversity metrics such as Shannon and Simpson indices as well as spatial distribution patterns. This approach reduces reliance on costly field inventories, remains robust across geographic contexts, and provides a scalable pathway for integrating biodiversity insights into urban resilience planning.
@@ 2 Results
We evaluated our framework in eight North American cities spanning a wide range of forest compositions and genera imbalances (see Section 4.1). The method consistently recovered meaningful biodiversity patterns, demonstrating robustness even in cities dominated by a few genera. Table 1 reports results for the best-performing configuration, selected through ablation and hyperparameter tuning (see Appendix B and 4), which achieved a balanced trade-off between richness accuracy, clustering quality, and diversity preservation.
@@
As shown in Table 1, genera imbalance influenced richness estimation error, but not uniformly across cities. In highly uneven forests such as Washington and Seattle, $RMSE_{\alpha}$ exceeded the within-city variability, confirming that richness estimation remains challenging under strong dominance patterns. In contrast, cities with more balanced compositions, including Los Angeles, Denver, and New York, achieved errors smaller than or comparable to their observed standard deviations, even improving substantially over a baseline mean predictor (See Appendix C for details). Despite variability, the V-score remained fairly consistent across all datasets (0.397–0.498), indicating that clustering preserved the overall partitioning structure even when richness estimates deviated. These results highlight that while richness is not the primary strength of our approach, performance is competitive in balanced systems and improves upon naïve baselines.
@@
In our datasets, Simpson diversity is bounded between 0 and 1, while Shannon diversity can reach values of about 3.5 (see Section 4.1). To enable direct comparison across indices and cities, Shannon values were normalized by their relative maximum, placing both metrics on a common [0,1] scale. In this context, Wasserstein distances below 0.05 indicate very strong alignment with ground truth distributions, while values near 0.1 represent moderate but still ecologically meaningful agreement. Cities such as Washington (0.133 for Shannon, 0.033 for Simpson) and San Francisco (0.034 for Shannon, 0.028 for Simpson) demonstrated strong alignment for both indices, with proportionally smaller errors for Simpson. Fig. 1 illustrates this by comparing predicted and observed values in Los Angeles (Shannon) and New York (Simpson). A full set of city-level results, including baseline comparisons and variability relative to observed standard deviations, is provided in Appendix C, which contextualizes how performance varies across urban forest structures.
@@ 3 Discussion
From an applied perspective, the framework offers a scalable alternative to field-based inventories for monitoring urban tree diversity and can be integrated into urban forestry programs. Spatially explicit Shannon and Simpson maps at the city scale can guide planting strategies, highlight areas at risk of monocultures, and support temporal monitoring, while reducing the effort and cost of repeated ground surveys. Importantly, this directly addresses the central gap identified in the Introduction: the inability of existing methods to recover abundance distributions and spatial diversity patterns from unlabeled imagery. By overcoming this limitation, our approach provides a pathway for continuous, cost-effective biodiversity monitoring that can inform urban resilience planning.
@@ 5.5 Code availability
The code used for analysis is available at https://github.com/Releaf.
@@ Appendix B Ablation analysis
Table 5: Ablation study results. $RMSE_{\text{$\alpha$}}$ = RMSE of richness; $W_{\text{Shannon}}$ = Wasserstein distance for Shannon diversity.
@@
Appendix C Comparisons with Baselines
@@
To verify that our similarity-based elimination is indeed beneficial — and not equivalent to simply discarding samples at random — we compare it against a random elimination strategy with the same removal fraction. The comparison uses the V-score as a measure of cluster–ground truth agreement, computed after elimination. As shown in Table 9, similarity-based elimination consistently outperforms random elimination across all cities, with an average improvement of approximately 0.10 in V-score and a standard deviation of $\pm$0.02. This demonstrates that the method selectively removes genuinely inconsistent samples rather than degrading cluster quality through indiscriminate pruning.
@@
Appendix D Validation of extracted visual features

### v2

@@ Abstract
Urban tree biodiversity is critical for climate resilience, ecological stability, and livability in cities, yet most municipalities lack detailed knowledge of their canopies. Field-based inventories provide reliable estimates of Shannon and Simpson diversity but are costly and time-consuming, while supervised AI methods require labeled data that often fail to generalize across regions. We introduce an unsupervised clustering framework that integrates visual embeddings from street-level imagery with spatial planting patterns to estimate biodiversity without labels. Applied to eight North American cities, the method recovers genus-level diversity patterns with high fidelity, achieving low Wasserstein distances to ground truth for Shannon and Simpson indices and preserving spatial autocorrelation. This scalable, fine-grained approach enables biodiversity mapping in cities lacking detailed inventories and offers a pathway for continuous, low-cost monitoring to support equitable access to greenery and adaptive management of urban ecosystems.
@@
Biodiversity, Unsupervised Learning, Clustering, Streetviews
@@ 1 Introduction
Open access data sources, such as Street View Imagery (SVI), offer a scalable alternative for biodiversity monitoring due to their extensive spatial coverage [22]. However, SVI datasets are unlabeled, meaning the images do not specify the tree species present and require substantial pre-processing to extract ecological information suitable for diversity estimations. Identifying individual taxa from imagery is particularly challenging: trees of the same species may look very different depending on season, age, or management, while distinct species can appear deceptively similar. These difficulties make automated classification highly sensitive to subtle morphological cues. Current supervised ecological AI models [23, 24], while effective within their training domains, often underperform when applied to regions or taxa not well represented in training data, largely due to geographic domain shift [25].
Self-supervised learning offers a promising alternative by eliminating the need for labeled data [26]. However, when applied to limited or homogeneous datasets, these methods often under-perform due to overfitting and poor generalization, as robust representations typically require large and diverse training samples. Few clustering approaches have been explicitly designed to preserve abundance distributions in a way that enables accurate Shannon and Simpson estimation from unlabeled imagery. This capability is essential for producing ecologically meaningful biodiversity metrics from large-scale, unlabeled datasets.
In this work, we address the challenge of large-scale, label-free biodiversity monitoring by developing an unsupervised clustering approach that estimates genus-level urban tree diversity directly from street-level imagery. We apply the framework across eight North American cities with contrasting forest compositions and imbalance, and evaluate its ability to recover biodiversity metrics such as Shannon and Simpson indices as well as spatial distribution patterns. While our evaluation is necessarily constrained to cities with available ground-truth inventories, the results demonstrate that the approach generalizes across diverse urban forest contexts and provides a scalable pathway for integrating biodiversity insights into urban resilience planning.
@@ 2 Results
We evaluated our framework in eight North American cities spanning a wide range of forest compositions and genera imbalances (see Section 4.1). The method consistently recovered meaningful biodiversity patterns, demonstrating robustness even in cities dominated by a few genera. Table 1 reports results for the best-performing configuration, selected through hyperparameter tuning (see Appendix A) and ablation study (see Appendix B), which achieved a balanced trade-off between richness accuracy, clustering quality, and diversity preservation.
@@
As shown in Table 1, genera imbalance strongly influenced richness estimation error ($RMSE_{\alpha}$), though not uniformly across cities. In highly uneven forests such as Washington and Seattle, $RMSE_{\alpha}$ exceeded within-city variability, underscoring the difficulty of estimating richness under strong dominance patterns. By contrast, in more balanced systems such as Los Angeles, Denver, and New York, errors were smaller than or comparable to observed standard deviations, and in several cases outperformed a naïve mean-predictor baseline (see Appendix D). Despite these differences, the V-score remained consistent across all datasets (0.397–0.498), indicating that clustering preserved the overall partitioning structure even when richness estimates deviated. These findings suggest that richness is not the primary strength of our approach, but performance is competitive in balanced forests and improves over statistical baselines.
@@
In our datasets, Simpson diversity is bounded between 0 and 1, while Shannon diversity can reach values of about 3.5 (see Section 4.1). To enable direct comparison across indices and cities, Shannon values were normalized by their relative maximum, placing both metrics on a common [0,1] scale. In this context, Wasserstein distances below 0.05 indicate very strong alignment with ground truth distributions, while values near 0.1 represent moderate but still ecologically meaningful agreement. Cities such as Washington (0.133 for Shannon, 0.033 for Simpson) and San Francisco (0.034 for Shannon, 0.028 for Simpson) demonstrated strong alignment for both indices, with proportionally smaller errors for Simpson. Calgary, by contrast, exhibited the weakest alignment across metrics. We hypothesize that this may be linked to more aggressive visual obfuscation apparent in Canadian Street View imagery compared to the United States, where stronger privacy safeguards and stricter blurring protocols are enforced [29]. Fig. 1 illustrates this by comparing predicted and observed values in Los Angeles (Shannon) and New York (Simpson). A full set of city-level results, including baseline comparisons and variability relative to observed standard deviations, is provided in Appendix D, which contextualizes how performance varies across urban forest structures.
Diagnosis of possible entropy inflation.
One potential concern with unsupervised clustering is that lower distributional errors may arise not from genuine recovery of community structure but from artificial entropy inflation, e.g., through oversplitting or the proliferation of singleton clusters. To assess this, we performed a dedicated diagnostic analysis (Appendix C) that examined oversplitting factors, purity, evenness shifts, and rank–abundance correlations between predicted and observed communities. Across all cities, the singleton ratio was zero, oversplitting factors remained close to one, and evenness shifts were modest, with New York even showing a slight decrease. At the same time, rank–abundance correlations were consistently high ($\rho\approx 0.9$), indicating that predicted clusters preserved the overall shape of community distributions. These findings confirm that the improved alignment of Shannon and Simpson diversity under the unsupervised approach is not an artifact of inflated entropy, but rather reflects ecologically meaningful recovery of biodiversity structure.
@@ 3 Discussion
From an applied perspective, the framework offers a scalable alternative to field-based inventories for monitoring urban tree diversity and can be integrated into urban forestry programs. Spatially explicit Shannon and Simpson maps at the city scale can guide planting strategies, highlight areas at risk of monocultures, and support temporal monitoring, while reducing the effort and cost of repeated ground surveys. Importantly, this directly addresses the central gap identified in the Introduction in regards to the inability of existing methods to recover abundance distributions and spatial diversity patterns from unlabeled imagery. By overcoming this limitation, our approach provides a pathway for continuous, cost-effective biodiversity monitoring that can inform urban resilience planning.
@@ 5.5 Code availability
The code used for analysis is available at https://github.com/Diaa340/Releaf/.
@@ Appendix B Ablation analysis
Table 5: Ablation study results.
@@
Appendix C Supervised Diagnostics
Table 6: Linear probing results (supervised baseline). Reported are RMSE for species richness and Wasserstein distances ($W_{1}$) for Shannon and Simpson indices.
City $RMSE_{\alpha}$ $\uparrow$ $W_{1}$–Shannon $\downarrow$ $W_{1}$–Simpson $\downarrow$
Calgary 1.91 0.306 0.140
New York 2.62 0.283 0.114
Columbus 3.25 0.405 0.143
Denver 3.40 0.320 0.093
Los Angeles 6.75 0.644 0.157
Seattle 2.47 0.413 0.132
Washington, D.C. 3.48 0.299 0.092
San Francisco 6.82 0.361 0.062
Linear probing provides a strong supervised baseline, yielding low richness RMSE across most cities. However, Wasserstein distances for Shannon and Simpson distributions remain relatively high, particularly in more diverse settings such as Los Angeles and San Francisco. This indicates that while supervised embeddings are precise for richness prediction, they do not fully capture the community structure of urban biodiversity.
Tables 7 and 8 report diagnostic indicators for the unsupervised clustering pipeline. Were the reduced Wasserstein distances merely the result of entropy inflation, we would expect pronounced oversplitting, the proliferation of singleton clusters, substantial positive shifts in evenness, and weak rank–abundance correlations. Instead, across all cities the singleton ratio was zero, oversplitting factors remained close to one, shifts in evenness were small (including a slight decrease in New York), and rank–abundance correlations were consistently high ($\rho\approx 0.9$). These results indicate that the unsupervised approach preserves the ecological structure of community distributions, rather than achieving lower errors through artificial entropy inflation.
[…]
Calgary 1.32 0.744 0.098
New York 0.70 0.582 0.027-0.027
Columbus 1.47 0.713 0.141
Denver 1.14 0.610 0.036
Los Angeles 1.03 0.641 0.076
Seattle 1.73 0.661 0.103
Washington, D.C. 1.55 0.578 0.046
San Francisco 1.13 0.555 0.039
Table 8: Distributional fidelity of the unsupervised pipeline. High $\rho$ values and modest Shannon/Simpson gaps indicate preservation of rank–abundance structure rather than entropy inflation.
City $\rho$ (Rank Shape $\uparrow$) Shannon Gap $\downarrow$ Simpson Gap $\downarrow$
Calgary 0.754 0.508 0.133
New York 0.928 0.486 0.153
Columbus 0.840 0.667 0.192
Denver 0.932 0.376 0.078
Los Angeles 0.913 0.438 0.112
Seattle 0.900 0.643 0.124
Washington, D.C. 0.912 0.567 0.069
San Francisco 0.940 0.404 0.071
Appendix D Comparisons with Statistical Baselines
@@
To verify that our similarity-based elimination is indeed beneficial, and not equivalent to simply discarding samples at random, we compare it against a random elimination strategy with the same removal fraction. The comparison uses the V-score as a measure of cluster–ground truth agreement, computed after elimination. As shown in Table 12, similarity-based elimination consistently outperforms random elimination across all cities, with an average improvement of approximately 0.10 in V-score and a standard deviation of $\pm$0.02. This demonstrates that the method selectively removes genuinely inconsistent samples rather than degrading cluster quality through indiscriminate pruning.
@@
Appendix E Validation of extracted visual features